OpenStudy (anonymous):

Given the geometric sequence where a1 = 4 and the common ratio is 3, what is the domain for n? All integers where n ≥ 1 All integers where > 1 All integers where n ≥ 4 All real number

OpenStudy (mosaic):

"n" is the number of terms. So it is a positive integer. The least it can be is 1.

OpenStudy (anonymous):

but is it this > or the one with line underneath

OpenStudy (mosaic):

1 is included. So \(n \ge 1\)

OpenStudy (jdoe0001):

\(\large r={\color{brown}{ 3}}\qquad \begin{array}{ccllll} term&value \\\hline\\ a_1&4\cdot {\color{brown}{ 3}}\\ a_2& (4\cdot 3)\cdot {\color{brown}{ 3}}\\ a_3&(4\cdot 3\cdot 3)\cdot {\color{brown}{ 3}}\\ a_4&(4\cdot 3\cdot 3\cdot 3)\cdot {\color{brown}{ 3}}\\ ....& ....\\ a_n& \square ? \end{array}\) where do you think that may end ? what could be "n"? and also notice it starts off from 4 onwards

OpenStudy (jdoe0001):

anyhow... as you can see, the geometric sequence starts off from 4 onwards and onwards to \(+\infty\)

OpenStudy (anonymous):

so n > with line under 4 then?

OpenStudy (anonymous):

Yes that would be correct